OpenStudy (anonymous):

A gift box is 12 inches long, 8 inches wide, and 2 inches high. How much wrapping paper is needed to wrap the gift box? A)272 in2 B)192 in2 C)384 in2 D)184 in2

OpenStudy (shane_b):

All you need to do is calculate the area of each side and add them together. If you need a nudge with that let me know.

jimthompson5910 (jim_thompson5910):

You can also use the formula SA = 2*(LW + WH + LH) where SA = surface area L = length W = width H = height

OpenStudy (anonymous):

i am sorry what ???

jimthompson5910 (jim_thompson5910):

where are you stuck?

jimthompson5910 (jim_thompson5910):

the formula for the surface area of a box with length L, width W and height H is SA = 2*(LW + WH + LH)

jimthompson5910 (jim_thompson5910):

in your case the length is 12, so L = 12 the width is 8, so W = 8 the height is 2, so H = 2

jimthompson5910 (jim_thompson5910):

which means SA = 2*(LW + WH + LH) turns into SA = 2*(12*8 + 8*2 + 12*2) after you plug the given info in

OpenStudy (anonymous):

ok so i have to multiply???

jimthompson5910 (jim_thompson5910):

yes, then add this is all for the stuff in the parenthesis

jimthompson5910 (jim_thompson5910):

once you have evaluated the stuff in the parenthesis, you just double the result

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

tell me what you get